25% Part (a) What net force is acting on her if she is standing on the merry-go-round 1.15 m from its center in newtons? The net force in this situation is sometimes called centripetal force. Grade Summary Deductions Potential Late Work 70% sino cos0 tano 7 8 9 Late Potential 70% 4 5 6 cotan0 asino acos0 Submissions 1 2 3 Attempts remaining: 10 at ano acotan0 sinh00 s per detailed view cosho tanho cotanho ODegrees Radians Submit Hint I give up! Hints: 2 deduction per hint. Hints remaining: 2 Feedback: 2 deduction per feedback. 25% Part (b) What net force (in N) s acting on her if she is standing on an amusement park merry-go-round that rotates at 2.2 rpm and she is 9.3 m from its center? 25% Part (d) How many times her weight is the force in part (b)?
Answer #1

Part a mass child mニ 19.75 K 叩㎡=.kr= y1-1.oey.mu n 6 0 6 0 15 N 4203N 60 60 949 N 그-42073 = 2-1구 193.74 0 05 I 13.74

answered by: ANURANJAN SARSAM
Answer #2

Part a mass child mニ 19.75 K 叩㎡=.kr= y1-1.oey.mu n 6 0 6 0 15 N 4203N 60 60 949 N 그-42073 = 2-1구 193.74 0 05 I 13.74